How can i solve this problem?
A ferry leaves a harbor and travels to a resort island at an average speed of 14mph. On the return trip, because of fog, the ferry travels at an average speed of 10mph. The total time for the trip is 6h. How far is the island from the harbor?

Answer by richwmiller(17219)

a+b=6,
14a=10b
a = 5/2=2.5, b = 7/2=3.5
35 miles=3.5*10